# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ Abstract
The escalating influx of data generated by networked edge devices, coupled with the growing awareness of data privacy, has promoted a transformative shift in computing paradigms from centralized data processing to privacy-preserved distributed data processing. Federated analytics (FA) is an emerging technique to support collaborative data analytics among diverse data owners without centralizing the raw data. Despite the wide applications of FA in industry and academia, a comprehensive examination of existing research efforts in FA has been notably absent. This survey aims to bridge this gap by first providing an overview of FA, elucidating key concepts, and discussing its relationship with similar concepts. We then conduct a thorough examination of FA, including its taxonomy, key challenges, and enabling techniques. Diverse FA applications, including statistical metrics, set computation, frequency-related applications, database query operations, model-based applications, FL-assisting FA tasks, and other wireless network applications are then carefully reviewed. We complete the survey with several open research issues and future directions. This survey intends to provide a holistic understanding of the emerging FA techniques and foster the continued evolution of privacy-preserving distributed data processing in the emerging networked society.
@@ I Introduction
There’s been a surge in data volume generated recently by the exponential growth of the Internet of Things (IoT) devices, where the number of connected devices is forecasted to reach 125 billion by 2030[1]. This IoT expansion is expected to generate 79.4 ZB of data by 2025[2]. Uploading and processing these data centrally present huge challenges for data communication and computing. Coincident with the extraordinary increase in the volume of data has also been a growing appreciation of the importance of data privacy, reflected by the enactment of stringent regulations such as the General Data Protection Regulation (GDPR) [3]. Fueled by the data explosion and privacy concerns, there is a noticeable shift from centralized to distributed data collection, storage, and processing. This evolution in data management underscores a new paradigm called federated computing, where data-oriented tasks are conducted among distributed data owners without uploading the local raw data. Federated learning (FL) is a instance of this computing paradigm that targets at collaborative model training without centralizing the raw data. It has been widely studied, deployed in real world, and adequately surveyed[4][5][6].
@@
Federated analytics (FA) is another emerging instance of the federated computation paradigm, which conducts data analytics tasks on distributed data held by different data owners (clients). The term “federated analytics” was first introduced by Google in 2020 [7] to describe “the practice of applying data science methods to the analysis of raw data that is stored locally on users’ devices”. The first application of FA is conducted by Google researchers [7], where the FA scheme is designed to evaluate the accuracy of a trained FL model against client data in local devices. FA extends the application of currently successful federated learning (FL) to broader data analytics tasks other than model training for deep learning. In contrast to classical data analytics workflow, which usually requires the clients to upload their raw data to a centralized server, FA prevents any transmission of raw data leaving the client they originated, in order to save communication cost and preserve the data privacy of clients.
@@
In an FA algorithm, the clients utilize their local data and the computation model received from the server to conduct local computation. The output of the local computation procedure is called “insight”. The insights are tailored data structures or information that reflect useful knowledge about the local data for the host data analytics task while preserving the privacy of the raw data. The clients upload the insight to the server to circumvent the raw data transmission, and the server performs insight aggregation to transform the individually-generated insights received from the clients to derive the population-level data analytics results. With FA, valuable analytics results can be derived to serve the pervasive data analytics needs of networked applications. The growing data privacy concerns of the users can also be properly handled.
Specifically, compared with traditional centralized data analytics and its sibling federated learning technique, FA offers the following benefits:
@@
  • Communication reduction: FA transmits the insight of client data instead of the raw data. The structure of insight is usually designed to contain only valuable information to resolve the data analytics tasks, and the size of insight is usually independent of the number of data samples held by the clients. Therefore, FA can reduce the communication cost in data analytics, especially when the clients hold large-volume raw data with many data samples.
@@
  • Privacy preservation: The tenet of FA prevents the transmission of raw data. The transmitted insight, which only includes abstractive information of the raw data, is less sensitive than the raw data transmission. Therefore, compared to traditional data analytics that require raw data transmission, FA enhances privacy preservation in utilizing edge data, with both the prevention of raw data transmission and additional privatization mechanisms. FA can be applied to both large-scale mobile devices, as well as several major data silos.
@@
  • Expansive task coverage: Different from FL which mostly focuses on neural network-based predictive tasks, FA extends the scope and covers the whole spectrum of descriptive tasks in data science. These tasks range from basic calculations, such as computing averages, to more complex operations in fields like graph analytics or video analytics. FA, joining forces with FL, successfully addresses the entirety of the learning and analytics problem space, enabling all data-oriented tasks to be completed in a federated environment.
@@
With these unique advantages, FA has offered a range of applications across various industries and sectors. For example, Google has utilized FA to improve federated learning models for its Gboard application [7] that it collects data from users’ keyboard activities to test the accuracy of its next-word prediction model. FA can also assist FL in terms of client selection [8], client clustering [9], and improving the privacy of aggregation [10]. Furthermore, sectors where data sensitivity is paramount can benefit from privacy-preserving FA to cooperate on tasks and research without sharing raw data such as healthcare[11], finance, and public services[12]. In these contexts, even competitive parties can conduct federated analytics collaboratively for common interests. The reductions in communication and computation further enable FA to offer wide applications in sectors where data is collected from remote locations like agriculture and environmental monitoring or where large-scale data collections and processing are required such as tasks in smart cities and in industry 4.0. The wide applications and studies of FA make it the proper time to survey this prominent area.
I-A Comparision and our contributions
Although the extraordinary advantage of the FA framework in privacy preservation and high-utility data analytics has spurred numerous FA studies, a systematic review of the current status of FA has not been conducted thus far. To date, only two preliminary surveys on FA exist[13, 14]. Survey [13] first reviews the federated analytics, and clarifies its position in the research literature, motivation, and application of it. However, it lacks detailed elaborations on tasks that can be supported by FA, presenting only two cases as examples. They place more emphasis on discussing the challenges, existing and possible solutions of FA in terms of the architecture, privacy problem, computation and communication resource, analytics design, and business models. Authors in [14] focus on common queries of interest in FA and their corresponding existing solutions, algorithms, and applications. They summarize and divide queries in FA into three categories: statistical queries, set-based queries, and matrix transformation queries. However, this classification is not comprehensive enough to cover other query types and data analytics tasks in FA, such as database operations, model-based queries, and tasks for FL assistance. Additionally, they do not provide a comprehensive summary of common techniques employed in FA, such as privatization, analytics, and system optimization. To the best of our knowledge, there is still no existing work to provide a comprehensive and dedicated review of federated analytics.
This survey is motivated by this gap to provide a more comprehensive summarization of FA algorithms and systems. We investigate wider classes of data analytics tasks FA applies than previous surveys. We clarify the ambiguity and discuss the overlap between FA and its related fields, showing how research in related fields exactly utilizes the idea of FA to resolve data analytics tasks with privacy preservation, and include these massive works in this survey. This survey is expected to let the readers have a comprehensive view of what the data analytics problem FA studies focus on, what are the enabling technologies of FA, and what are the opening issues of FA that require further research effort. In addition, this survey is expected to inspire researchers in related fields about the relationship and application of FA in these fields. It provides a demonstration of how FA has been applied in these fields, and how FA can be potentially utilized in these fields.
To this end, the key contributions of this article are highlighted as follows:
  • We present the first state-of-the-art comprehensive survey on FA, covering the taxonomy, key challenges, and techniques, as well as a wide spectrum of FA applications.
  • We derive a taxonomy for FA from different dimensions, including data analytics tasks, scale of clients, number of iterations, coordination model, as well as threat models.
  • We summarize enabling techniques in FA for privacy preservation, data analytics, and system optimization.
  • We summarize a wide spectrum of FA applications, including statistical metrics, set computation, frequency-related applications, database operations, model-based applications, other downstream and upstream tasks in a machine learning pipeline, and wireless network applications.
  • We identify several key open research issues and challenges in FA, covering the application scenarios, algorithm design, system optimization, and cross-layer enhancement.
@@ I-B Structure of the Survey
The structure of the survey is organized as follows. Section II provides an overview of FA, including its general workflow, a general mathematical problem formulation, and its comparison with related topics. Section III introduces our taxonomy of FA algorithms and systems. Our taxonomy covers five important dimensions that have significant impacts on its real-world applications and deployments. In Section IV, key challenges in designing and deploying FA are discussed. Section V, we introduce key enabling techniques applied in FA, including privatization techniques to realize privacy preservation, analytics techniques to derive data analytics results, and deployment/optimization techniques to enable real-world applications. In the following five sections, we introduce the vast existing FA solutions in detail, where each section covers FA solutions on a particular class of data analytics tasks: Section VI covers the computation of simple statistical metrics; Section VII covers the tasks related to frequency estimation; Section VIII covers database operations (data analytics tasks formulated as SQL queries); Section IX covers the diverse model-based data analytics problems; Section X covers data analytics tasks that are tailor-designed to assist FL systems; Section XI covers the applications of FA for wireless networks. Section XII discusses open issues and future research directions in FA, and Section XIII concludes the survey. An overview of the survey is provided in Fig. 1.
Figure 2: Architecture of FA.
II Overview of Federated Analytics

### v3

@@ Abstract
The escalating influx of data generated by networked edge devices, coupled with the growing awareness of data privacy, has restricted the traditional data analytics workflow, where the edge data are gathered by a centralized server to be further utilized by data analysts. To continue leveraging vast edge data to support various data-incentive applications, computing paradigms have promoted a transformative shift from centralized data processing to privacy-preserved distributed data processing. The need to perform data analytics on private edge data motivates federated analytics (FA), an emerging technique to support collaborative data analytics among diverse data owners without centralizing the raw data. Despite the wide applications of FA in industry and academia, a comprehensive examination of existing research efforts in FA has been notably absent. This survey aims to bridge this gap by first providing an overview of FA, elucidating key concepts, and discussing its relationship with similar concepts. We then thoroughly examine FA, including its key challenges, taxonomy, and enabling techniques. Diverse FA applications, including statistical metrics, frequency-related applications, database query operations, FL-assisting FA tasks, and other wireless network applications are then carefully reviewed. We complete the survey with several open research issues, future directions, and a comprehensive lessons learned part. This survey intends to provide a holistic understanding of the emerging FA techniques and foster the continued evolution of privacy-preserving distributed data processing in the emerging networked society.
@@ I Introduction
There’s been a surge in data volume generated recently by the exponential growth of the Internet of Things (IoT) devices, where the number of connected devices is forecasted to reach 125 billion by 2030[1]. This IoT expansion is expected to generate 79.4 ZB of data by 2025[2]. The utilization of these data drives numerous significant and predominant data-driven applications in the fields of science and industry, ranging from chemical design to recommender systems. As the utilization of these edge data increases, uploading and processing these data centrally presents huge challenges for data communication and computing. Coincident with the extraordinary increase in the volume of data has also been a growing appreciation of the importance of data privacy, reflected by the enactment of stringent regulations such as the General Data Protection Regulation (GDPR) in EU [3] and the California Consumer Privacy Act (CCPA) in California [4]. Fueled by the data explosion and privacy concerns, there is a noticeable shift from centralized to distributed data collection, storage, and processing. This evolution in data management underscores a new paradigm called federated computation (also known as federated computing) [5, 6, 7], where data-oriented tasks are conducted among distributed data owners without uploading the local raw data. In federated computation, distributed data owners convert their local data into informative and privacy-preserving intermediate outcomes (such as gradients in federated learning), and upload them to a centralized server. The intermediate outcome is the transformed result based on certain models specified by the designer from the local dataset (e.g., gradient derived from backward propagating ML models specified by the designer). The data analyst, who controls the centralized server, aggregates these intermediate outcomes to derive a global model, or analytics results, which represents a global view of the edge data in all data owners.
@@
While federated computation serves as the foundational infrastructure for distributed computation, federated analytics, and federated learning are distinct but closely related concepts that operate within this framework. Federated learning (FL) is an instance of this computing paradigm that targets collaborative model training without centralizing the raw data. It has been widely studied, deployed in the real world, and adequately surveyed [8, 9, 10, 11, 12]. Federated analytics (FA) is another emerging instance of the federated computation paradigm, which conducts data analytics tasks on distributed data held by different data owners (clients). The term “federated analytics” was first introduced by Google in 2020 [13] to describe “the practice of applying data science methods to the analysis of raw data that is stored locally on users’ devices”. FA extends the application of currently successful FL to broader data analytics tasks other than model training for deep learning. For instance, the first application of FA is conducted by Google researchers [13], where the FA scheme is designed to evaluate the accuracy of a trained FL model against client data in local devices. The primary distinction between FA and FL lies in their different focused tasks, in that FA focuses on descriptive tasks, and FL focuses on predictive tasks. For instance, while FL focuses on training a machine learning model across multiple devices, such as in predictive text algorithms on smartphones, FA can be applied to tasks like aggregating statistical summaries such as collecting statistics, frequent pattern mining, location heatmap, and healthcare data analysis. These tasks do not involve procedures of machine learning model training, so FL cannot be applied. In contrast to classical data analytics workflow, which usually requires the clients to upload their raw data to a centralized server, FA prevents any transmission of raw data leaving the client they originated, with in order to save communication costs and preserve the data privacy of clients.
@@
In an FA algorithm, the clients utilize their local data and the computation model received from the server to conduct local computation. The output of the local computation procedure is called “insight” in this paper and it is one form of “intermediate outcome” in federated computation as we discussed previously. The insights are tailored data structures or information that reflect useful knowledge about the local data for the host data analytics task while preserving the privacy of the raw data. The clients upload the insight to the server to circumvent the raw data transmission, and the server performs insight aggregation to transform the individually generated insights received from the clients to derive the population-level data analytics results. With FA, valuable analytics results can be derived to serve the pervasive data analytics needs of networked applications. The growing data privacy concerns of the users can also be properly handled.
Specifically, FA offers the following benefits:
@@
  • Communication reduction: FA, following the paradigm of federated computation, transmits the insight of client data instead of the raw data. The structure of insight is usually designed to contain only valuable information to resolve the data analytics tasks, and the size of insight is usually independent of the number of data samples held by the clients. Therefore, FA can reduce the communication cost in data analytics, especially when the clients hold large-volume raw data with many data samples.
@@
  • Privacy preservation: The tenet of federated computation lets FA prevent the transmission of raw data. The transmitted insight, which only includes abstractive information of the raw data, is less sensitive than the raw data transmission. Therefore, compared to traditional data analytics that requires raw data transmission, FA enhances privacy preservation in utilizing edge data, with both the prevention of raw data transmission and additional privatization mechanisms. FA can be applied to both large-scale mobile devices, as well as several major data silos.
@@
  • Expansive task coverage: Different from FL which mostly focuses on neural network-based predictive tasks, FA extends the scope and covers the whole spectrum of descriptive tasks in data science. These tasks range from basic calculations to complex tasks involving sophisticated and dedicated data structures and computation procedures. FA, joining forces with FL, successfully addresses the entirety of the learning and analytics problem space, enabling all data-oriented tasks to be completed in a federated environment.
@@
With these unique advantages, FA has offered a range of applications across various industries and sectors. For example, Google has utilized FA “in support of FL” to “measure the quality of FL models against real-world data when that data is not available in a data center” [13]. FA can also assist FL in terms of client selection [14] and client clustering [15]. Furthermore, sectors where data sensitivity is paramount can benefit from privacy-preserving FA to cooperate on tasks and research without sharing raw data such as healthcare[16], finance, and public services[17]. In these contexts, even competitive parties can conduct FA collaboratively for common interests. The reductions in communication and computation further enable FA to offer wide applications in sectors where data is collected from remote locations like agriculture and environmental monitoring or where large-scale data collections and processing are required such as tasks in smart cities and in industry 4.0. The wide applications and studies of FA make it the proper time to survey this prominent area.
I-A Comparison and our contributions
While numerous surveys have explored related topics such as FL [11, 10, 9, 8, 12] and privacy computing [18, 19, 20], a comprehensive review of the current status of FA is still lacking. Only two preliminary surveys exist. Survey [21] offers an overview of FA, its research positioning, motivations, and applications but lacks in-depth task elaboration, focusing more on challenges and solutions. Survey [22] categorizes FA queries into statistical, set-based, and matrix transformation tasks, but fails to address the full range of data analytics tasks in contemporary FA applications with significant performance implications. Additionally, it overlooks a multi-dimensional taxonomy and a thorough analysis of key enabling techniques from various perspectives.
In comparison to these previous surveys [21, 22], this article provides a more comprehensive summary of FA algorithms and systems. We explore a broader set of data analytics tasks to which FA is applied, clarifying ambiguities and examining the overlap between FA and related fields. Our analysis highlights how research in these adjacent areas utilizes FA concepts to address data analytics challenges while preserving privacy. We also integrate these extensive works into this survey. The goal of this article is to offer readers a well-rounded view of the data analytics problems that FA addresses, the enabling technologies behind FA, and the open research questions that require further exploration. Furthermore, we aim to inspire researchers in related fields by demonstrating the relationship between FA and its applications, showcasing how FA has been successfully implemented, and suggesting how it could be further leveraged in these domains.
The key contributions of this paper are outlined as follows:
  • Comprehensive survey. We provide a state-of-the-art, comprehensive survey on FA, covering key challenges, taxonomies, techniques, and a wide range of FA applications. It offers a holistic view of FA by addressing not only its core principles but also the critical gaps that have persisted in existing literature.
  • Taxonomy development. We present a robust multi-dimensional taxonomy for FA, encompassing aspects such as data analytics tasks, client scale, iteration count, coordination models, and privatization methodologies. It not only provides a comprehensive framework for analyzing existing FA systems but also serves as a guide for designing future solutions that align with the specific needs of diverse applications.
  • Enabling techniques. We summarize the enabling techniques in FA that address privacy preservation, data analytics, and system optimization. They collectively address the dual challenges of maintaining high data utility and strong privacy guarantees, while ensuring that FA systems are scalable, resource-efficient, and practical for real-world applications.
  • Application spectrum. We compile an extensive range of FA applications, including statistical metrics, frequency-related tasks, database operations, machine learning pipeline tasks, and wireless network applications. These applications demonstrate how FA addresses key challenges in privacy-sensitive, large-scale, and distributed environments, making it a transformative technology for industries ranging from IoT to cloud computing and beyond.
  • Open research issues. We identify several critical open research challenges in FA, spanning application scenarios, algorithm design, system optimization, and cross-layer enhancements. These open research challenges outline a roadmap for future exploration and highlight the potential for FA to evolve into a more versatile and impactful framework for privacy-preserving distributed analytics.
TABLE I: List of important abbreviations, ordered by their occurrence in this survey.
FA Federated analytics
IoT Internet of things
FL Federated learning
DP Differential privacy
MEC Mobile edge computing
MPC Multi-party computation
DDM Distributed data mining
LDP Local differential privacy
CDP Central differential privacy
DDP Distributed differential privacy
HE Homomorphic encryption
PSI Private set intersection
FPM Frequent pattern mining
SVM Support vector machine
@@ I-B Structure of the survey
The structure of the survey is organized as follows. Section II reviews surveys of FA and related topics. Section III provides an overview of FA, including its general workflow, a general mathematical problem formulation, and its comparison with related topics. It also includes our taxonomy of FA works that covers five important dimensions that have significant impacts on its real-world applications and deployments. In Section IV, key challenges in designing and deploying FA are discussed. Section V, we introduce key enabling techniques applied in FA, including privatization techniques to realize privacy preservation, analytics techniques to derive data analytics results, and deployment/optimization techniques to enable real-world applications. In the following five sections, we introduce the vast existing FA solutions in detail, where each section covers FA solutions on a particular class of data analytics tasks: Section VI covers the computation of simple statistical metrics; Section VII covers the tasks related to frequency estimation; Section VIII covers database operations (data analytics tasks formulated as SQL queries); Section IX covers data analytics tasks that are tailor-designed to assist FL systems; Section X covers the applications of FA for wireless networks. Section XI discusses open issues and future research directions in FA, and Section XII concludes the survey. An overview of the survey is provided in Fig. 1. For reading convenience, we summarize the important abbreviations used in this survey and conclude them in Table I.
Figure 2: Architecture of FA. The left plot demonstrates the interaction between the server and the clients. The clients holding privacy-sensitive local data utilizes the computation procedure of insight extractor to derive privacy-preserving insight, which is aggregated on the server side to derive results. The right plot demonstrates the iterative workflow of FA, where an FA iteration consists of phases of computation model distribution, local computation, insight upload, and insight aggregation.
II Related Surveys
In this section, we begin by first conducting a discussion of the existing survey on FA [21, 22], systematically reviewing its current research landscape and development trends. Subsequently, we extend our focus to include surveys on two closely related topics: FL [11, 10, 9, 8, 12] and Privacy Computing [18, 19, 20]. These surveys offer valuable insights into the foundational technologies, methodologies, and challenges closely tied to FA, providing a broader perspective on the research and development of this domain.
II-A Federated analytics survey
Although the extraordinary advantage of the FA framework in privacy preservation and high-utility data analytics has spurred numerous FA studies, a systematic review of the current status of FA has not been conducted thus far. To date, only two preliminary surveys on FA exist[21, 22]. Survey [21] first reviews the FA, and clarifies its position in the research literature, motivation, and application of it. However, it lacks detailed elaborations on tasks that can be supported by FA, presenting only two cases as examples. They place more emphasis on discussing the challenges, existing and possible solutions of FA in terms of the architecture, privacy problem, computation and communication resource, analytics design, and business models. Authors in [22] focus on common queries of interest in FA and their corresponding existing solutions, algorithms, and applications. They summarize and divide queries in FA into three categories: statistical queries, set-based queries, and matrix transformation queries. However, this classification is not comprehensive enough to cover other query types and data analytics tasks in FA, such as database operations, frequency-based queries, and tasks for FL assistance. Additionally, they do not provide a comprehensive summary of common techniques employed in FA, such as privatization, analytics, and system optimization.
II-B Federated learning survey
FL has obtained significant attention in recent years as a promising approach to distributed machine learning, allowing multiple devices or parties to collaboratively train models without sharing raw data. This paradigm shift addresses critical issues related to data privacy, security, and communication costs, making it particularly relevant in various applications such as healthcare, finance, and IoT. Numerous surveys [11, 10, 9, 8, 12] have been conducted to explore the advancements, challenges, and future directions of FL.
The survey [11] provides an in-depth analysis of recent advances in FL and its future directions. It summarizes the fundamental principles of FL and its advantages over traditional centralized learning approaches and reviews the latest developments in FL, including advancements in communication efficiency, personalization techniques, and secure aggregation methods. They discuss how these innovations have been applied in various sectors, such as finance, healthcare, and autonomous systems, showcasing the versatility and effectiveness of FL. In survey [10], authors comprehensively discuss the challenges, methods, and future directions of FL. It begins by outlining the primary issues in FL, such as data privacy, communication costs, and model accuracy. They review various methods to address these challenges, including differential privacy (DP), secure multiparty computation, and model compression techniques. They also examine the application of FL in various fields like healthcare, finance, and IoT, highlighting the benefits and limitations of FL in these contexts.
There are FL surveys concentrating on specific area. The survey [9] focuses on the application of FL in mobile edge networks . It discusses the unique challenges of implementing FL in such environments, including heterogeneous devices, communication costs, and privacy concerns. The authors provide an overview of mobile edge computing (MEC) and its role in facilitating FL, detailing how FL can optimize resource allocation and improve network efficiency. They review existing solutions for integrating FL with MEC, such as edge aggregation and adaptive model updating. The survey also explores various use cases of FL in mobile networks, such as vehicular networks and smart cities. While survey [8] examines the use of FL in IoT domain. It introduces the key characteristics of IoT, such as vast data generation and the need for real-time processing, which make FL an attractive solution. The authors review the main challenges in applying FL to IoT, including data heterogeneity, device resource constraints, and security issues. They present various strategies to overcome these challenges, such as lightweight models, federated transfer learning, and privacy-preserving techniques. The survey also highlights several applications of FL in IoT, such as smart homes, industrial IoT, and healthcare monitoring. Also in [12], researchers focus on the application of FL in the IoT domain, highlighting recent advancements, a comprehensive taxonomy, and open challenges. They introduce the necessity of FL in IoT due to the massive data generated by distributed devices and the privacy concerns associated with centralized learning and then review recent advancements in FL techniques tailored for IoT, emphasizing metrics such as sparsification, robustness, quantization, scalability, security, and privacy. The survey also presents a detailed taxonomy of FL approaches in IoT, categorizing them based on various parameters such as optimization schemes, incentive mechanisms, security and privacy measures, and aggregation methods.
II-C Privacy computing survey
With the proliferation of data-driven applications and the increasing awareness of privacy concerns, privacy-preserving computing has become a critical area of research. Various surveys have been conducted to explore the different aspects of privacy computing, particularly in the context of edge computing [18, 19, 20].
In [18], the authors explore the privacy and security issues inherent in edge computing. It details the architecture of edge computing and how it differs from traditional cloud computing. The authors discuss the primary security challenges such as data confidentiality, integrity, and availability, highlighting the unique threats posed by the distributed nature of edge environments. They provide an overview of existing security solutions, including encryption techniques, authentication protocols, and access control mechanisms. The survey also examines privacy-preserving methods, emphasizing the need for lightweight and efficient solutions to handle the resource-constrained nature of edge devices. Similarly, the survey [19] focuses on the security and privacy aspects of multi-access edge computing. It begins by outlining the key components and functionalities and explaining its importance in enhancing computational capabilities at the network edge. The authors review various security challenges specific to multi-access edge computing, such as secure data transmission, user authentication, and intrusion detection. They highlight current solutions and strategies employed to mitigate these issues, including blockchain technology, secure multi-party computation (MPC), and machine learning-based anomaly detection. The survey also addresses privacy concerns, particularly data anonymization and DP techniques, to protect user information in multi-access edge computing environments. In [20], it also delves into data security and privacy-preserving mechanisms within the edge computing paradigm. It discusses the benefits of edge computing over centralized cloud models, particularly in terms of reducing latency and improving data locality. The authors examine the specific security and privacy challenges faced by edge computing, such as secure data storage, secure computation, and user authentication. They provide a comprehensive review of cryptographic techniques like homomorphic encryption, attribute-based encryption, and proxy re-encryption that can be applied to secure data in edge environments. Additionally, the survey covers access control mechanisms and privacy-preserving data aggregation methods.
III Overview of Federated Analytics